Two employees of a local flower shop, Pat an fulfill online FA) and d Jen, are both qualified to make flower arrangements ( ennn(O0). Their production possibilitics are below, showing combinations of flower arrangements and fulfilled online orders an employee can complete in one hour 12 Pat Jea (3,6) 10 00 What is the opportunity cost of fulfilling 1 online order in terms of flower arrangements for Pat? What is the opportunity of fulfilling 1 online order in terms of flower arrangements for Jan? Calculate. Who has a comparative advantage in fulfilling online orders? a) Using information from the graphs, can you assign the two employees to a specialized task of either making flower arrangements or fulfilling online orders? b) Originally, Pat and Jen were both doing flower arrangements and online orders. Pat was able to accomplish 3 О.О. and 6 F.A. in an hour, while Jen did 5 O.O. and 4 F.A. How would specialization affect their productivity? Calculate and compare their joint hourly output before and after they specialize. c) (Bonus) Suppose the shops demand has dropped, and now every hour it sells 9 flower arrangements, and needs to fulfill 5 online orders. The shop owner is thinking about laying off Jen, while sending Pat to get super training in processing online orders. How could Pats productivity (and PPF) change in order to be able to fulfill shops demand? Draw Pats new PPF, indicating both vertical and horizontal intercepts What skill should the training emphasize? Should it emphasize one or both skills? (Write on reverse if needed) d)
Answer #1

a) The opportunity cost of fulfilling 1 online order in terms of flower arrangement for Pat is 12/6 =2 and for Jen the opportunity cost of fulfilling 1 online order in terms of flower arrangement is 8/10 = 0.8. As 0.8< 2 , so Jen has comparative advantage in fulfilling 1online order.

b) From the information given in the graph we can say that Jen has a specialisation in online order because of comparative advantage and more production in one hour in terms of online order. We can also say that similarly Pat has comparative advantage in flower arrangement because Pat has to sacrifice less online order to do an extra flower arrangement. Because Pat has to sacrifice 6/12=0.5 online for one extra flower arrangement but Jen has to sacrifice 10/8 =1.25 online order for extra flower arrangement. So Pat has specialisation in flower arrangement.

c) Before specialisation Pat was doing 3 online order and 6 flower arrangement in an hour and Jen was doing 5 online order and 4 flower arrangement in an hour. So in total they were doing 8 online (5+3) order and 10(6+4) flower arrangement in an hour. But after specialisation Pat will do 12 flower arrangement and Jen will be capable to do 10 online order. So after specialisation they are able to 10-8 =2 extra online order and 12-10 =2 extra flower arrangement. Their joint output by taking online order and flower arrangement will be 10(online order)+12(flower arrangement) =22 after specialisation but before specialisation they both able to make 8(online order)+ 10(flower arrangement) =18.

d) If shop's demand falls to 9 flower arrangement and 5 online in an hour and if the shop owner send Pat in training in online order and after training Pat's productivity in online order will increase and it's PPF will shift rightward. The new PPF will be as follows- 1 4 12 10上 0yder

So there should give training on both online order and flower arrangement so that Pat can able to do 5 online order and 9 flower arrangement in an hour. There should emphasize on both.